OpenStudy (anonymous):

t^2-13.125t I need the quadratic expression please help

OpenStudy (anonymous):

i'm realy confused as to what to do from here

OpenStudy (anonymous):

the equation u gave is a quadratic expression. what should we do with it/. should we find roots of it or wat?

OpenStudy (anonymous):

i'm trying to find the amnt of time a jet of water is in the air here is the full equation 0=-16t^2+210t+0

OpenStudy (anonymous):

Do u know how to solve to a quadratic equation?

OpenStudy (anonymous):

not this particular one :(

OpenStudy (anonymous):

Can u solve x^2-5x+6 =0 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I got this far x^2=5x-6

OpenStudy (anonymous):

but i think that's wrong

OpenStudy (anonymous):

for a quadratic equation \[ax^2+bx+c =0\] roots or factoors are \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

i gave you that eqution x^2 -5x+6 =0 to know whether u can solve quadratic equations or not! dont worry!

OpenStudy (anonymous):

ok thank you for all ur help I suck at math

OpenStudy (anonymous):

solve for \[16t^2-210t+0=0\] solve for t from the formula that ive given, you will get t values 0 and some other value, that value of t which is not zero is your requird answer!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

t = 0 and t = 13s so 13seconds is the answer